OpenStudy (anonymous):

x varies inversely as y2, and x = 4 when y = 10. Find x when y = 2. A. x = 16 B. x = 5 C. x = 80 D. x = 100

OpenStudy (anonymous):

doryy cant give you just the answer

OpenStudy (anonymous):

Can you help me work it out then?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

do you recall the 2 basic forms that this type of question takes?

OpenStudy (anonymous):

No.

OpenStudy (amistre64):

the is the direct variation: y = kx and the indirect variation: y = k/x which one of these do you spose this question would need to use

OpenStudy (anonymous):

I don't know, honestly... I haven't paid attention to math since 7th grade and now I'm in 12th and I greatly regret it.

OpenStudy (amistre64):

just humor me and take a guess, its a 50/50 shot

OpenStudy (anonymous):

y = k/x

OpenStudy (amistre64):

lol, very good :)

OpenStudy (amistre64):

so, lets use this form as it pertains to the question

OpenStudy (amistre64):

"x varies inversely as y^2" tells us: y^2 = k/x "x = 4 when y = 10" are the specifics 10^2 = k/4 cam you solve for k? we need it for the next step

OpenStudy (anonymous):

uh 20?

OpenStudy (amistre64):

... not even close, but nice try ;)

OpenStudy (anonymous):

Haha

OpenStudy (amistre64):

10^2 = k/4 100 = k/4 4(100) = k 400 = k

OpenStudy (anonymous):

oh okay

OpenStudy (amistre64):

now, lets delete the x,y values and keep the k y^2 = 100/x ; and it asks us to find x when y = 2, so lets plug it in 2^2 = 100/x how would you suggest we solve for "x"?

OpenStudy (anonymous):

uhhh i have no idea >.< I probably seem like a dumbass. :\

OpenStudy (amistre64):

and k=400, not 100 ... but thats just these fat little fingers trying to type :)

OpenStudy (amistre64):

these are basic algebra techniques that you will certainly have to practice more .... 2^2 = 400/x 4 = 400/x 4x = 400 x = 400/4 now what would you say x is equal to?

OpenStudy (anonymous):

100

OpenStudy (amistre64):

yay!!, i agree ;)

OpenStudy (anonymous):

so that's the answer then right?

OpenStudy (amistre64):

i hope so, or else im gonna have to go back to 3rd grade ....

OpenStudy (anonymous):

>.< alright haha thanks

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

How would I do this one? http://prntscr.com/kvs6b

OpenStudy (amistre64):

that one would take a lot of work that you presently do not have the background to solve ... post it as a new question, and let someone else have the chance to get a medal by guiding you thru it

OpenStudy (anonymous):

alright. :)